OpenStudy (anonymous):

MEDAL!! Solve equation, giving a general formula for all the solutions (cos x - sqrt2/2)(sec x-1)=0

OpenStudy (anonymous):

I know the answers are x=2pi, 0, pi/4, and 7pi/4. I don't know the general formula

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

well general formula is always 2kpi + min angle

OpenStudy (anonymous):

ok… lol I still don't know

OpenStudy (ikram002p):

u have x =2pi right ? general formula means all angle are congruent to 2 pi so 2pi +2pi is congruent to 2pi 2pi +4pi is congruent to 2pi 2pi +6pi is congruent to 2pi .... so the general formula 2pi +2kpi is congruent to 2pi 

OpenStudy (anonymous):

2pi + 2kpi is the general formula? are there more formulas or would that be it?

OpenStudy (ikram002p):

so u have x=0 , general formula 0 + 2kpi x=pi/4 , general formula pi/4+2kpi x=7pi /4 , general formula 7pi/4 +2kpi

OpenStudy (anonymous):

Oh ok, so I have 4 general formulas for this question thanks!

OpenStudy (ikram002p):

np , ur wlc !

OpenStudy (anonymous):

ALso, for these it is 2kpi but for the last question it was just koi. correct?

OpenStudy (ikram002p):

nope its always 2kpi , last one i made a typo check it again

OpenStudy (anonymous):

are you 100% sure it is ALWAYS 2kpi?

OpenStudy (ikram002p):

yep ^_^

OpenStudy (anonymous):

ok thanks!

OpenStudy (ikram002p):

np , ur wlc again !